OpenStudy (anonymous):

What is the solution set of |2x – 3| – 6 = –1? 5 or –5 4 or 1 4 or –1 No solution

OpenStudy (rc94):

4 or -1

OpenStudy (nurali):

|2x – 3| – 6 = –1 |2x – 3| = –1+6 |2x – 3| = 5 2x – 3=5 or 2x – 3=-5 2x =5+3 or 2x =-5+3 x=4 or x=-1 so the answer is 4 or –1